Can someone please help asap??The government conducts a survey of the population and finds that 115 of 664 respondents report voting for party A last elect
Answer #1

p1cap = X1/N1 = 115/664 = 0.1732
p1cap = X2/N2 = 104/715 = 0.1455
pcap = (X1 + X2)/(N1 + N2) = (115+104)/(664+715) = 0.1588

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: p1 = p2
Alternate Hypothesis, Ha: p1 ≠ p2

Рі — Р2 Z= УР(1 - Р)(1/п +1/п2)
Test statistic
z = (p1cap - p2cap)/sqrt(pcap * (1-pcap) * (1/N1 + 1/N2))
z = (0.1732-0.1455)/sqrt(0.1588*(1-0.1588)*(1/664 + 1/715))
z = 1.41